# RG theory of spontaneous stochasticity for Sabra model of turbulence

### v1

@@ IV.2 RG relation between the flows
The first two maps are related to the space-scaling transformation (13). Together with the projection $\pi_{1}$ they yields the identity as $\sigma_{+}\circ\sigma_{-}=\pi_{1}+\sigma_{-}\circ\sigma_{+}=\mathrm{id}$. For each map (32), we denote the respective (Dirac) probability kernel by the capital latter $\Sigma_{+}$, $\Sigma_{-}$ and $\Pi_{1}$; see Appendix VIII.3. We also define the map
@@
$\Phi^{(N+1)}_{T,\mathbf{b}}(\mathbf{s}_{\mathrm{new}}|\mathbf{s})=\left[\Pi_{1}\ast\Sigma_{-}\circ\big{(}\Phi^{(N)}_{t^{\prime},\mathbf{b}^{\prime}}\big{)}^{\infty}\circ\Sigma_{+}\right]\left(\mathbf{s}_{\mathrm{new}}|\mathbf{s}^{\prime}\right),\quad\mathbf{s}^{\prime}=\xi_{T,\mathbf{b}}(\mathbf{s}),\quad\mathbf{b}^{\prime}=2(b_{0},u^{\prime}_{1}).$ (35)
@@
Figure 7: Using the self-similar structure of DT-LES model with the cutoff $N+1$, we represent the dynamics at shells $n\geq 2$ in terms of the flow corresponding to the cutoff $N$.
Consider the random map $\varphi_{T,\mathbf{b}}^{(N+1)}$ from Eq. (27) for the increased cutoff $N+1$. This map contains the algorithm steps (21), (22) and (24) between two consecutive synchronized states; see Fig. 4. The first step corresponds to $n=1$ and, hence, yields the state $\mathbf{s}^{\prime}$ given by the map (33); recall that the viscous and noise step is not present because $N+1\geq 3$; see Eq. (26). The remaining steps affect only the components $n\geq 2$ and, therefore, can be represented by a random map $\psi$ acting on the shifted state $\sigma_{+}(\mathbf{s}^{\prime})$; see Fig. 7. This yields $\mathbf{s}_{\mathrm{new}}=\pi_{1}(\mathbf{s}^{\prime})+\sigma_{-}\circ\psi\circ\sigma_{+}(\mathbf{s}^{\prime})$, where $\sigma_{-}$ is the backward shift and $\pi_{1}$ restores the first component. For the probability kernels, this yields
@@
where $\Psi$ is the probability kernel corresponding to the random map $\psi$, and the sum is replaced by a convolution; see Appendix VIII.3. Finally, as a consequence of the scaling symmetry (18), one can see that the map $\psi$ is given by essentially the same algorithm; see Fig. 7. This algorithm is formulated for the new final time $t^{\prime}$, the shifted boundary state $\mathbf{b}^{\prime}$, the cutoff $N$, and also the loop steps (25) must be skipped. Thus, similarly to the relation (29) we obtain $\Psi=\big{(}\Phi_{t^{\prime},\mathbf{b}^{\prime}}^{(N)}\big{)}^{\infty}$ with the same meaning that an infinite sequence of compositions converges in a finite number steps. Combining this with Eq. (36) proves the theorem. ∎
The RG relations of Theorem 1 extend to a wide class of DT-LES models. One can verify that the same relations are valid for any dissipative and noise terms whose dependence on the cutoff NN has scaling symmetry (18). In such models, the flows for different $N$ are related by the same RG operator (34).
@@ V.1 Ideal limit as a dynamical system
The main consequence of the RG relation (34) is that it represents the ideal limit of $N\to\infty$ as a dynamical system in the space of flows. The RG operator $\mathcal{R}$, which determines this dynamics, depends only on the ideal system, i.e., on the nonlinear interactions $B_{1}^{+}$ and $B_{1}^{-}$ determining the map (33) through Eq. (22) with $n=1$. This operator does not depend on the specific form of dissipation and noise at small scales, in particular, on the values of the coefficients $D$ and $C$ in case (26). Which particular small-scale regularization is used is “encoded” in the flow $\Phi^{(N)}$ and is passed on to the next flow $\Phi^{(N+1)}$ due to scaling symmetry (18). In other words, the ideal model determines the RG operator, while the choice of regularization determines the initial flow $\Phi^{(2)}$ for the RG dynamics.
By representing the ideal limit as the RG dynamics, we also identify the probability distributions obtained in this limit with an attractor of this dynamics. We now formulate the concept of a fixed-point RG attractor, which is based on the assumption that the classical theory of dynamical systems is applicable; see e.g. [1]. Of course, this does not follow automatically, since the RG dynamics occurs in an infinite-dimensional space of flows, but our numerical results will strongly support this assumption. The existence and stability of the RG attractor provides a theoretical explanation for spontaneous stochasticity, including its universality property and convergence process.
@@ V.2 Fixed-point RG attractor
A flow $\Phi^{\infty}$ is the fixed-point RG attractor if it satisfies the fixed-point relation $\mathcal{R}[\Phi^{\infty}]=\Phi^{\infty}$ and has the convergence property
@@
Here $\mathcal{B}(\Phi^{\infty})$ denotes a basin of attraction in a space of flows. Using relation (31), we write the ideal limit for the family of transition probabilities as
@@
These limits apply to the DT-LES models if the respective flows belong to the basin of attraction.
@@
We can now relate the existence of the RG attractor to the four properties of spontaneous stochasticity listed Section II.1. ($i$) The RG attractor implies naturally the convergence for respective probability distributions. ($ii$) The stochasticity property follows if the limiting family of transition probabilities $\mathrm{P}^{\infty}$ is not Dirac. (iiiiii) The spontaneity property requires that realizations of the limiting process solve the ideal model. This is true because the dynamics at each given scale is governed by ideal couplings for $N$ sufficiently large. ($iv$) Finally, universality with respect to the formulation of the ideal limit is a consequence of the attracting condition (38). This universality refers to all types of small-scale dissipation and noise, for which the respective flows: (a) satisfy the RG relation $\Phi^{(N+1)}=\mathcal{R}[\Phi^{(N)}]$ and (b) belong to the basin of attraction.
@@ V.3 Linearization and stability of a fixed-point RG attractor
where $\delta\mathcal{R}$ is a linear operator acting on flow perturbations $\Psi$. Note that $\Psi$ denotes a family of kernels $\Psi_{T,\mathbf{b}}(\mathbf{s}_{\mathrm{new}}|\mathbf{s})$, which are signed measures with respect to $\mathbf{s}_{\mathrm{new}}$. Substituting Eq. (41) into the RG relation (34) and denoting $\Psi^{(N)}=\Phi^{(N)}-\Phi^{\infty}$, we obtain the linearized dynamics
@@
The stability condition is formulated as $|\rho|<1$, in which case the correction (43) decays as $N\to\infty$. It follows that both $\rho$ and $\Omega$ are universal, i.e. they do not depend on the regularization parameters (small-scale dissipation and noise). The prefactor $c\in\mathbb{R}$ is the only part of the expression (43) that is not non-universal.
@@
where the second expression integrates the observable $O$ with respect to the eigenvector kernel $\mathrm{Q}_{T,\mathbf{b}}$ similarly to Eq. (39). Recall that the RG operator and, hence, the eigenmode depend on the ideal system only. Therefore, the quantities in the asymptotic relation (48) have different forms of universality: (a) The eigenvalue $\rho$ is fully universal in the sense that it does not depend on the initial and boundary conditions, the final time, small-scale regularization, or the observable. (b) The integral $q$ depends on the initial and boundary conditions, the final time and on the observable, but is independent of the regularization. (c) Finally, the prefactor $c$ depends on the regularization, but does not depend on the initial and boundary conditions, the final time, and the observable. Recall that the independence of regularization is assumed here within the class of models to which the RG theory is applicable and whose flows belong to the basin of the RG attractor.
@@ V.4 Numerical verification of the universal RG eigenvalue
Here the error bounds corresponding to variations in different measurements. Solid lines in Fig. 8 show the right-hand sides of Eq. (49) with the eigenvalue (50) and optimized values for $C$ and $\phi$. In all cases, the measured ensemble averages show very good agreement with the RG prediction for large $N$. It is noteworthy that $|\rho|$ in Eq. (50) is close to unity, which explains the slow convergence in the ideal limit.
@@
Finally, we verify numerically the universality of the RG eigenvalue. We do this by considering a different regularization with different initial and boundary conditions. First, instead of the large-scale initial state $\mathbf{a}=(1,0,0,\ldots)$, we consider the “rough” initial condition in the form of geometric progression $\mathbf{a}=(z,z^{2},z^{3},\ldots)$, where $z=2^{h}\exp(3i)$ with the Hölder exponent $h=0.3$. We use the same boundary conditions as before with the final time $T=1$. A typical dynamics is presented in Fig. 9(a). The moments $\big{\langle}|u_{1}(T)|\big{\rangle}$ and $\big{\langle}|u_{1}(T)|^{2}\big{\rangle}$ computed for different $N$ are shown in Figs. 9(b,c). The solid lines present the asymptotic relations (49), where we used the same RG eigenvalue (50) and tuned prefactors and phases.
Figure 9: (a) Dynamics of the LES model for rough initial data. (b,c) Dependence on the moments $\big{\langle}|u_{1}(T)|\big{\rangle}$ and $\big{\langle}|u_{1}(T)|^{2}\big{\rangle}$ on NN (empty dots) fitted with the universal RG eigenmode prediction (solid lines). Statistics used $10^{5}$ simulations, and statistical errors are comparable with the dot size. (d-f) Similar results for the different small-scale regularization (51).
@@ VI RG approach to the fluctuating Sabra model
The RG approach developed in the previous sections was based on the key property that regularized models preserve the scaling symmetries of the ideal model: the time scaling symmetry (12) and the extended form of the space scaling symmetry (18). Two examples are the LES models given by Eqs. (17) and (51). We call such symmetry-preserving regularizations canonical. It is the self-similarity of the dissipative and noise terms in the canonical regularization that leads to the existence of the RG operator $\mathcal{R}$ and, thus, to a representation of the ideal limit $N\to\infty$ by the RG dynamics of flows $\Phi^{(N)}$.
@@
with a fixed integer parameter $M$. The dissipation and noise terms in Eq. (52) can be seen as characterized by the effective wave number $k_{*}$ and the turnover time $\tau_{*}$. The regularizations (52) are canonical, i.e. they possess both symmetries (12) and (18). One can check that discrete-time versions of models (52) yield the same RG operator from Theorem 1: if $\Phi^{(N)}$ is the flow for the model (52), then $\Phi^{(N+1)}=\mathcal{R}[\Phi^{(N)}]$ is the flow for the increased cutoff shell $N+1$. Strictly speaking, this derivation requires neglecting the dissipative and noise terms at shell $n=1$, which is valid in the inviscid limit $N\to\infty$. Hence, the ideal limit $N\to\infty$ for each auxiliary model (52) is spontaneously stochastic and universal. We confirm this numerically in Fig. 3 for the parameters $D=1$, $C=0.01$, $M=4$ and $N=24$.
@@
When $\mathrm{Re}^{-1}$ and $\sqrt{\Theta}$ are small, for example in the ideal limit formulated in (5), one can choose the time-dependent $M(t)$ such that both $D(t)\lesssim 1$ and $C(t)\lesssim 1$. This is illustrated in Fig. 10 for a specific realization of the Sabra model with $\mathrm{Re}=10^{-9}$ from Fig. 1(a). The correspondence (54) means that, at every time instant, the Sabra model behaves as the auxiliary model with specifically chosen parameters. The RG theory predicts that the dynamics of the latter is approximated by the universal spontaneously stochastic process. Therefore, the same process approximates the dynamics of the fluctuating Sabra model.

### v2

@@ IV.2 RG relation between the flow kernels
The first two maps are related to the space-scaling transformation (13). Together with the projection $\pi_{1}$ they yields the identity as $\sigma_{+}\circ\sigma_{-}=\pi_{1}+\sigma_{-}\circ\sigma_{+}=\mathrm{id}$. For each map (33), we denote the respective (Dirac) probability kernel by the capital latter $\Sigma_{+}$, $\Sigma_{-}$ and $\Pi_{1}$; see Appendix VIII.4. We also define the map
@@
$\Phi^{(N+1)}_{T,\mathbf{b}}(\mathbf{s}_{\mathrm{new}}|\mathbf{s})=\left[\Pi_{1}\ast\Sigma_{-}\circ\big(\Phi^{(N)}_{t^{\prime},\mathbf{b}^{\prime}}\big)^{\mathrm{fin}}\circ\Sigma_{+}\right]\left(\mathbf{s}_{\mathrm{new}}|\mathbf{s}^{\prime}\right),\quad\mathbf{s}^{\prime}=\xi_{T,\mathbf{b}}(\mathbf{s}),\quad\mathbf{b}^{\prime}=2(b_{0},u^{\prime}_{1}).$ (36)
As before, $\big(\Phi^{(N)}_{t^{\prime},\mathbf{b}^{\prime}}\big)^{\mathrm{fin}}$ denotes the composition $\Phi^{(N)}_{t^{\prime},\mathbf{b}^{\prime}}\circ\cdots\circ\Phi^{(N)}_{t^{\prime},\mathbf{b}^{\prime}}$, where the number of kernels is sufficient to reach the final time $t^{\prime}$.
@@
Figure 7: Using the self-similar structure of DT-LES model with the cutoff $N+1$, we represent the dynamics at shells $n\geq 2$ in terms of the flow kernel corresponding to the cutoff $N$.
Consider the kernel $\Phi_{T,\mathbf{b}}^{(N+1)}(\mathbf{s}_{\mathrm{new}}|\mathbf{s})$ from Eq. (27) for the increased cutoff $N+1$. It describes the probability transition $\mathbf{s}\mapsto\mathbf{s}_{\mathrm{new}}$ corresponding to the algorithm steps (21), (22) and (24); see Fig. 4. The first step, which corresponds to $n=1$, yields the state $\mathbf{s}^{\prime}$ given by the map (34); recall that the viscous and noise step is not present because $N+1\geq 3$; see Eq. (26). The remaining steps affect only the components $n\geq 2$ and, therefore, can be represented by a kernel $\Psi$ describing the transition of the shifted state $\sigma_{+}(\mathbf{s}^{\prime})\mapsto\mathbf{s}^{\prime\prime}$; see Fig. 7. This yields $\mathbf{s}_{\mathrm{new}}=\pi_{1}(\mathbf{s}^{\prime})+\sigma_{-}(\mathbf{s}^{\prime\prime})$, where $\sigma_{-}$ is the backward shift and $\pi_{1}$ restores the first component. For the probability kernels, this yields
@@
where the sum of states is replaced by a convolution of kernels; see Appendix VIII.4. Finally, as a consequence of the scaling symmetry (18), one can see that the kernel $\Psi$ is given by essentially the same algorithm; see Fig. 7. This algorithm is formulated for the new final time $t^{\prime}$, the shifted boundary state $\mathbf{b}^{\prime}$, the cutoff $N$, and also the loop steps (25) must be skipped. Thus, similarly to the relation (30) we obtain $\Psi=\big(\Phi_{t^{\prime},\mathbf{b}^{\prime}}^{(N)}\big)^{\mathrm{fin}}$. Combining this with Eq. (37) proves the theorem. ∎
The RG relations of Theorem 1 extend to a wide class of DT-LES models. One can verify that the same relations are valid for any dissipative and noise terms whose dependence on the cutoff $N$ has scaling symmetry (18). In such models, the flow kernels for different $N$ are related by the same RG operator (35).
@@ V.1 Ideal limit as a dynamical system
The main consequence of the RG relation (35) is that it represents the ideal limit of $N\to\infty$ as a dynamical system in the space of flow kernels. The RG operator $\mathcal{R}$, which determines this dynamics, depends only on the ideal system, i.e., on the nonlinear interactions $B_{1}^{+}$ and $B_{1}^{-}$ determining the map (34) through Eq. (22) with $n=1$. This operator does not depend on the specific form of dissipation and noise at small scales, in particular, on the values of the coefficients $D$ and $C$ in case (26). Which particular small-scale regularization is used is “encoded” in the flow kernel $\Phi^{(N)}$ and is passed on to the next flow kernel $\Phi^{(N+1)}$ due to scaling symmetry (18). In other words, the ideal model determines the RG operator, while the choice of regularization determines the initial flow kernel $\Phi^{(N_{0})}$ for the RG dynamics.
By representing the ideal limit as the RG dynamics, we also identify the probability distributions obtained in this limit with an attractor of this dynamics. We now formulate the concept of a fixed-point RG attractor, which is based on the assumption that the classical theory of dynamical systems is applicable; see e.g. [1]. Of course, this does not follow automatically, since the RG dynamics occurs in an infinite-dimensional space of flow kernels, but our numerical results will strongly support this assumption. The existence and stability of the RG attractor provides a theoretical explanation for spontaneous stochasticity, including its universality property and convergence process.
@@ V.2 Fixed-point RG attractor
A flow kernel $\Phi^{\infty}$ is the fixed-point RG attractor if it satisfies the fixed-point relation $\mathcal{R}[\Phi^{\infty}]=\Phi^{\infty}$ and has the convergence property
@@
Here $\mathcal{B}(\Phi^{\infty})$ denotes the basin of attraction in the space of flow kernels, meaning that $\mathcal{B}(\Phi^{\infty})$ contains an open neighborhood of the fixed point $\Phi^{\infty}$ in an appropriate mathematical sense. Using relation (32), we write the ideal limit for the family of transition probabilities as
@@
These limits apply to the DT-LES models if the respective flow kernels belong to the basin of attraction.
@@
We can now relate the existence of the RG attractor to the four properties of spontaneous stochasticity listed Section II.1. ($i$) The RG attractor implies naturally the convergence for respective probability distributions. ($ii$) The stochasticity property follows if the limiting family of transition probabilities $\mathrm{P}^{\infty}$ is not Dirac. (iiiiii) The spontaneity property requires that realizations of the limiting process solve the ideal model. This is true because the dynamics at each given scale is governed by ideal couplings for $N$ sufficiently large. ($iv$) Finally, universality with respect to the formulation of the ideal limit is a consequence of the attracting condition (39). This universality refers to all types of small-scale dissipation and noise, for which the respective flow kernels: (a) satisfy the RG relation $\Phi^{(N+1)}=\mathcal{R}[\Phi^{(N)}]$ and (b) belong to the basin of attraction.
We remark that the RG formalism can be introduced at the level of the continuous-time model, similarly to [36], as long as the model can be formulated as a finite system of SDEs. This condition, however, does not hold at the fixed point, and therefore the RG operator is not well defined there. Consequently, the linearized RG problem developed below also loses a clear mathematical meaning. These difficulties are resolved at the level of the discrete-time model.
@@ V.3 Linearization and stability of a fixed-point RG attractor
where $\delta\mathcal{R}$ is a linear operator acting on perturbations $\Psi$ of the flow kernel. Note that $\Psi$ denotes a family of kernels $\Psi_{T,\mathbf{b}}(\mathbf{s}_{\mathrm{new}}|\mathbf{s})$, which are signed measures with respect to $\mathbf{s}_{\mathrm{new}}$. Substituting Eq. (42) into the RG relation (35) and denoting $\Psi^{(N)}=\Phi^{(N)}-\Phi^{\infty}$, we obtain the linearized dynamics
@@
The stability condition is formulated as $|\rho|<1$, in which case the correction (44) decays as $N\to\infty$. It follows that both $\rho$ and $\Omega$ are universal, i.e. they do not depend on the regularization parameters (small-scale dissipation and noise). The prefactor $c\in\mathbb{R}$ is the only part of the expression (44) that is not universal.
@@
where the second expression integrates the observable $O$ with respect to the eigenvector kernel $\mathrm{Q}_{T,\mathbf{b}}$ similarly to Eq. (40). Recall that the RG operator and, hence, the eigenmode depend on the ideal system only. Therefore, the quantities in the asymptotic relation (49) have different forms of universality: (a) The eigenvalue $\rho$ is fully universal in the sense that it does not depend on the initial and boundary conditions, the final time, small-scale regularization, or the observable. (b) The integral $q$ depends on the initial and boundary conditions, the final time and on the observable, but is independent of the regularization. (c) Finally, the prefactor $c$ depends on the regularization, but does not depend on the initial and boundary conditions, the final time, and the observable. Recall that the independence of regularization is assumed here within the class of models to which the RG theory is applicable and whose flow kernels belong to the basin of the RG attractor.
@@ V.4 Numerical verification of the universal RG eigenvalue
This fitting is obtained by minimizing a suitable norm of the difference between the left- and right-hand sides of Eq. (50) with respect to the unknown parameters $\langle O\rangle^{\infty}_{T,\mathbf{b},\mathbf{a}}$, $\rho$, $C$ and $\phi$. The error bounds in Eq. (51) correspond to variations across different measurements. Solid lines in Fig. 8 show the right-hand sides of Eq. (50) with the eigenvalue (51) and optimized values for $C$ and $\phi$. In all cases, the measured ensemble averages show very good agreement with the RG prediction for large $N$. It is noteworthy that $|\rho|$ in Eq. (51) is close to unity, which explains the slow convergence in the ideal limit.
@@
Finally, we verify numerically the universality of the RG eigenvalue. We do this by considering a different regularization with different initial and boundary conditions. First, instead of the large-scale initial state $\mathbf{a}=(1,0,0,\ldots)$, we consider the “rough” initial condition in the form of geometric progression $\mathbf{a}=(z,z^{2},z^{3},\ldots)$, where $z=2^{-h}\exp(3i)$ with the Hölder exponent $h=0.3$. We use the same boundary conditions as before with the final time $T=1$. A typical dynamics is presented in Fig. 9(a). The moments $\big\langle|u_{1}(T)|\big\rangle$ and $\big\langle|u_{1}(T)|^{2}\big\rangle$ computed for different $N$ are shown in Figs. 9(b,c). The solid lines present the asymptotic relations (50), where we used the same RG eigenvalue (51) and tuned prefactors and phases.
Figure 9: (a) Dynamics of the LES model for rough initial data. (b,c) Dependence on the moments $\big\langle|u_{1}(T)|\big\rangle$ and $\big\langle|u_{1}(T)|^{2}\big\rangle$ on NN (empty dots) fitted with the universal RG eigenmode prediction (solid lines). Statistics used $10^{5}$ simulations, and statistical errors are comparable with the dot size. (d-f) Similar results for the different small-scale regularization (52) for zero initial condition.
@@ VI RG approach to the fluctuating Sabra model
The RG approach developed in the previous sections was based on the key property that regularized models preserve the scaling symmetries of the ideal model: the time scaling symmetry (12) and the extended form of the space scaling symmetry (18). Two examples are the LES models given by Eqs. (17) and (52). We call such symmetry-preserving regularizations canonical. It is the self-similarity of the dissipative and noise terms in the canonical regularization that leads to the existence of the RG operator $\mathcal{R}$ and, thus, to a representation of the ideal limit $N\to\infty$ by the RG dynamics of flow kernels $\Phi^{(N)}$.
@@
with a fixed integer parameter $M$. The dissipation and noise terms in Eq. (53) can be seen as characterized by the effective wave number $k_{*}$ and the turnover time $\tau_{*}$. The regularizations (53) are canonical, i.e. they possess both symmetries (12) and (18). One can check that discrete-time versions of models (53) yield the same RG operator from Theorem 1: if $\Phi^{(N)}$ is the flow kernel for the model (53), then $\Phi^{(N+1)}=\mathcal{R}[\Phi^{(N)}]$ is the flow kernel for the increased cutoff shell $N+1$. Strictly speaking, this derivation requires neglecting the dissipative and noise terms at shell $n=1$, which is valid in the inviscid limit $N\to\infty$. Hence, the ideal limit $N\to\infty$ for each auxiliary model (53) is spontaneously stochastic and universal. We confirm this numerically in Fig. 3 for the parameters $D=1$, $C=0.01$, $M=4$ and $N=24$.
@@
When $\mathrm{Re}^{-1}$ and $\sqrt{\Theta}$ are small, for example in the ideal limit formulated in (5), one can choose the time-dependent $M(t)$ such that both $D(t)\lesssim 1$ and $C(t)\lesssim 1$. This is illustrated in Fig. 10 for a specific realization of the Sabra model with $\mathrm{Re}=10^{9}$ from Fig. 1(a). The correspondence (55) means that, at every time instant, the Sabra model behaves as the auxiliary model with specifically chosen parameters. The RG theory predicts that the dynamics of the latter is approximated by the universal spontaneously stochastic process. Therefore, the same process approximates the dynamics of the fluctuating Sabra model.